OpenStudy (anonymous):

solve equation (r-3)(r+6)=0

OpenStudy (anonymous):

it's already solved!! here r-3=0 or r+6=0 therefore the two values of r are 3 and -6

OpenStudy (mew55):

r^2+6r-18

OpenStudy (mew55):

i meant 3

OpenStudy (anonymous):

The question is asking which values of r will cause the left side to equal zero. Factored form is great because when 'any' of the factors equal zero then the left side equals zero. Knowing this you can set each of the two factors to zero and solve individually just as @ProgramGuru did: \[r-3=0\] \[r+6 =0\]

OpenStudy (anonymous):

o maybe i just didn't see it